40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6258     Entry time: Tue Feb 7 03:05:08 2012
Author: Zach 
Type: Update 
Category: SUS 
Subject: OSEM sat amp measurements 

I did some more investigation on the OSEM box today.

Troubleshooting:

After removing some capacitors and still finding that the +15V rail was at over +20V, I decided to see if the TO-3 7815 that I removed behaved properly all by itself. It did. After some more poking around, I discovered that whoever assembled the board isolated the case of the regulator from the board. It is through the case that this package gets its grounding, so I removed the mica insulator, remounted the regulator, and all worked fine.

Since I had gotten a spare from Downs, I also replaced the LT1031 (precision 10-V reference), for fear that it had been damaged by the floating voltage regulator.

 

Noise measurements:

sat_amp_measurement_setup.png

LED Driver

With the above out of the way, I was finally able to take some measurements. The first thing I did was to look at the LED drivers. I fixed the one stage that I mentioned in my last post by adding two 820-ohm resistors in parallel with the 1k, such that it was very close to all the others (which are 806 || 806 || 1k). With that, using a red LED, I measured a current of 34.5 mA (+/- 0.1) out of each of the 5 stages (UL, UR, LL, LR, S).

I then measured the current noise of each one by monitoring the voltage across the 287-ohm resistor in series with the LED. The driver works by putting the LED in the feedback path of an inverting amp. There is a 10-V input from the LT1031, and the values of the input and feedback resistors determine the current drawn through the LED. There is a buffer (LM6321) in the path to provide the necessary current.

The LISO model I made according to that description seems to make sense. I simply modeled the LED as a small resistor and asked LISO for the current through it. The transfer function shows the proper DC response of -49.15 dB(A/V) -->  34.8 mA @ 10 V, but, the estimated current noise doesn't add up with the measured levels:

LED_driver_vs_liso.png

I have to get to the bottom of this. Two possibilities are: 1) The buffer adds noise, and/or 2) I am modeling this invalidly.

PD Amp

I also began measuring the PD amplifier noise levels, though I only measured two of them for lack of time. I find it odd that there is a 100-ohm input series resistor on what I thought would be just a transimpedance amplifier. For that reason, I want to look into how the OSEMs are connected to this guy.

In any case, I measured the output noise of two of the PD amps by shorting the input side of the 100-ohm resistors to ground, and then I divided by their TF to get the input noise level. Here it is compared with the LISO estimate. I have plotted them in units of voltage noise at the input side of the resistors for lack of a way to infer the equivalent photocurrent noise level.

pd_amp_vs_liso.png

Above 2 Hz or so, the measured level agrees with the prediction. Below this, the measured noise level increases as 1/f, while it should go as the standard 1/sqrt(f) (the manufacturer-quoted 1/f corner is at 2 Hz). Another thing to get to the bottom of.

ELOG V3.1.3-